Nick is one-forth his grandfather's age. Five years ago, he was one-fifth his grandfather's age.

How old are Nick and his Grandfather now?

Answer by stanbon(75887)

Nick is one-forth his grandfather's age. Five years ago, he was one-fifth his grandfather's age.
-------------------
Equations:
n = (1/4)g
(n-5) = (1/5)(g-5)
---------
Substitute for "g" and solve for "n"::
n-5 = (1/5)(4n-5)
----
n-5 = (4/5)n - 1
-------
(1/5)n = 4
----
n = 20 yrs ( Nick's age now)
---
Solve for "g":
g = 4n = 80 years (Grandpa's age now)
--------------
Cheers,
Stan H.
----------